# DINAMO: Dynamic and INterpretable Anomaly MOnitoring for Large-Scale Particle Physics Experiments

### v1

@@ B.7 Full Pseudocode
B.8 Relations across metrics and robustness
Figure 7: Variation of the performance metrics for the DINAMO-S algorithm as a function of the value of the $\alpha$ parameter. Solid lines (bands) correspond to averages (standard deviations) over the ensemble of 1000 datasets.
The change in performance metrics for the DINAMO-S algorithm as a function of the value of $\alpha$ is shown in Figure 7, averaged over the 1000 datasets. The performance curves are found to be reasonably flat for a wide range of $\alpha$ values, which demonstrates the robustness of the method even if $\alpha$ is not optimally chosen for a given dataset.
@@ Appendix C DINAMO-ML Algorithm Details
In this appendix, we provide the complete technical description of our machine-learning approach, referred to as DINAMO-ML. While Section III.3 of the main text highlights its core concepts and motivations, here we elaborate on the Transformer encoder-based architecture, the way it constructs a conditional PDF of histogram bin contents, and the associated training/inference flow.
C.1 Method Overview
DINAMO-ML replaces the hand-crafted EWMA with a learnable transformation from the last $M$ good histograms to a predicted reference $\hat{\mu}$ and uncertainty $\hat{\sigma}$. Both $\hat{\mu}$ and $\hat{\sigma}$ are inferred via a Transformer encoder-based network that processes these $M$ previous good histograms (additionally, it takes M relative timestamps, where “relative” means “with respect to the current i-th one”). The algorithm continues to produce a $\chi^{2}$ and pull $\delta$, mirroring DINAMO-S in interpretability and overall workflow (predict-reference, compute-anomaly-score, decide good/bad, update), but it learns the temporal weighting instead of relying on a single smoothing factor $\alpha$. In practice, we set the anomaly threshold using historical data (e.g., a validation sample of labeled runs), optimizing for an acceptable trade-off between false positives and false negatives.
C.2 Transformer Encoder-based Reference Prediction
C.3 Training Procedure
Conditional PDF via Factorized Gaussian.
From a probabilistic standpoint, DINAMO-ML performs a mini-batch likelihood fit over good histograms. Concretely, for each good histogram $\bm{\tilde{x}}_{k}$, we treat each bin $\tilde{x}_{k,j}$ as a draw from a Gaussian:
$\tilde{x}_{k,j}\;\sim\;\mathcal{N}\Bigl{(}\hat{\mu}_{j},\;\hat{\sigma}_{j}^{2}\Bigr{)},$
where the $\hat{\mu}_{j}$ and $\hat{\sigma}_{j}$ (predicted by the neural network) depend on the conditional variables (the last $M$ good histograms), and $\bm{\tilde{x}}_{i}$ := $\bm{x}_{i}\ /\ I_{x_{i}}$, where $I_{x_{i}}$ is a sum across all bins $N_{b}$: $I_{x_{i}}=\sum_{j}^{N_{b}}{x_{i,j}}$. Minimizing the Gaussian negative log-likelihood (NLL) of this model forces the network to learn how best to map previous good data into the current reference parameters:
$NLL=-\!\log p\bigl{(}\bm{\tilde{X}}_{K}\bigl{)}=\frac{1}{2N_{b}K}\sum_{k=1}^{K}\sum_{j=1}^{N_{b}}\biggl{[}\left(\frac{\tilde{x}_{k,j}-\hat{\mu}_{k,j}}{\hat{\sigma}_{k,j}}\right)^{2}+\ \log{(\hat{\sigma}_{k,j}^{2})}\biggr{]},$
where $K$ is the mini-batch size, $N_{b}$ is the number of bins in the histograms, and ${\{\tilde{x}_{k}\}}_{k=1...K}$ are the normalized good histograms, each is taking by the model with its $M$ contextual previous good runs and their relative times.
Mini-Batch Updates.
To implement this in practice:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    We backpropagate to adjust the Transformer encoder’s parameters, effectively “learning” how to predict the bin-by-bin means and uncertainties for future runs.
  5. 5.
C.4 Inference and Anomaly Detection
  1. 1.
    Context Construction: For a new run $\bm{x}_{i}$, we identify up to $M$ most recent good histograms (with time $<$ $i$) to form a conditional context (again constant-padding if fewer than $M$ exist).
  2. 2.
    Predict Reference: We take the transformer encoder model outputs $\hat{\bm{\mu}}_{i},\hat{\bm{\sigma}}_{i}$.
  3. 3.
    Anomaly Score: Then, we compute
    $\chi^{2}_{\nu}\;=\;\frac{1}{N_{b}}\sum_{j=1}^{N_{b}}\frac{\bigl{(}\tilde{x}_{i,j}-\hat{\mu}_{i,j}\bigr{)}^{2}}{\sigma_{\tilde{x}_{i,j},p}^{2}+\hat{\sigma}_{\mu_{i,j}}^{2}}$
    and a pull vector $\bm{\delta}_{i}$:
    $\delta_{i,j}\;=\;\frac{\tilde{x}_{i,j}-\hat{\mu}_{i,j}}{\sqrt{\sigma_{\tilde{x}_{i,j},p}^{2}+\hat{\sigma}_{\mu_{i,j}}^{2}}}.$
    Here, $\bm{\tilde{x}}_{i}$ is run’s histogram normalized to unity, so that $\bm{\tilde{x}}_{i}=\bm{x}_{i}\ /\ I_{x_{i}}$ and $\bm{\sigma}_{\bm{\tilde{x}}_{i},p}=\sqrt{\bm{\tilde{x}}_{i}\ /\ I_{x_{i}}-\bm{\tilde{x}}^{2}_{i}\ /\ I_{x_{i}}}$ is the bin-wise estimation of the Poisson uncertainty for $\bm{\tilde{x}}_{i}$. To decide whether $\bm{x}_{i}$ is good or bad, the anomaly score is compared to the threshold.
  4. 4.
    Update: If $i$-th run is labeled good, we add $(\bm{\tilde{x}}_{i},i)$ to the dataset $\mathcal{D}$; if bad, we exclude it.
C.5 Full Pseudocode
Algorithm C.5 presents the full pseudocode of the DINAMO-ML method.
{algorithm}
[h] DINAMO-ML: Transformer Encoder-based Algorithm
1:  Input: $\{\bm{x}_{i},y_{i}\}_{i=1}^{N}$, buffer size $M$, batch size $K$, learning rate, etc.
2:  Initialize:
3:   $\mathcal{D}\leftarrow\varnothing$
@@ C.5 Full Pseudocode
5:   $\text{Optimizer}\leftarrow\textsc{AdamW}(\text{NNParams},\text{lr}=5\times 10^{-4},\text{weight\_decay}=10^{-4})$
6:   EarlyStopPatience $\leftarrow 5$
7:  for $i\leftarrow 1$ to $N$ do
8:     // 1. Normalize the new histogram
9:     Normalize $\bm{x}_{i}$ (if nonzero), get $\bm{\tilde{x}}_{i}$
10:     // 2. Perform inference to obtain reference prediction
11:     $\text{context}\leftarrow\textsc{BuildContext}(\mathcal{D},\,i,\,M)$
[…]
13:     // 3. Compute $\chi^{2}_{\nu}$ and compare to anomaly threshold to classify
14:     // 4. A posteriori label is assigned (by shifter or ground truth)
15:     if $y^{(i)}=0$ then
16:        $\mathcal{D}\leftarrow\mathcal{D}\cup(\bm{\tilde{x}}_{i},i)$
17:        // 5. Training step
18:        Select the $K$ most recent entries from $\mathcal{D}$ (if fewer than $K$ exist, take all).
19:        $nll_{\text{batch}}\leftarrow 0$
20:        Optimizer.zero_grad()
21:        for each $(\bm{\tilde{x}}_{k},i_{k})$ in the selected batch:
22:         $\text{context}\leftarrow\textsc{BuildContext}(\mathcal{D},\,i_{k},\,M)$
[…]
24:         $nll_{\mathrm{k}}\leftarrow\textsc{ComputeNLL}(\bm{\tilde{x}}_{k},\hat{\bm{\mu}},\hat{\bm{\sigma}})$
25:         $nll_{\text{batch}}\leftarrow nll_{\text{batch}}+nll_{\mathrm{k}}$
26:        $nll_{\text{batch}}\leftarrow nll_{\text{batch}}\;\big{/}\;\#(\text{selected batch})$
27:        Backprop on $nll_{\text{batch}}$
28:        CheckEarlyStopping(EarlyStopPatience)
29:     end if
30:     /* End of iteration for run $i$/
31:  end for
Model’s Hyperparameters.
Using a trial-and-error approach, the following effective hyperparameters were selected:
  • $M=20$ for the context buffer (the last 20 good runs).
  • $K=10$ for the mini-batch size.
[…]
  • Learning rate of $5\times 10^{-4}$ with the AdamW optimizer, weight decay of $10^{-4}$.
  • Early stopping with a patience of 5 based on the NLL loss.
C.6 Results on a Single Synthetic Dataset
In this section we provide results on a single synthetic dataset using DINAMO-ML model. We use the same dataset with seed of 70 that was introduced in the main text and previously used to assess the performance of the DINAMO-S model in Section IV.2.2.
Figure 8: Anomaly score of the DINAMO-ML algorithm as a function of the run number for a single dataset composed of 5000 runs. Red points represent the ground-truth bad runs and the blue points represent the good runs. The black dashed line is the optimized, using the historical regime, threshold to assign a class label: if a run is above it, the run is predicted to be bad and vice versa.
Figure 8 shows the logarithm of the reduced $\chi^{2}$ anomaly score as a function of run number. The first 1000 runs are used in so-called “historical” or “validation” regime for the hyperoptimisation tune. The dashed black line represent the optimized, using the historical data, separation threshold. In contrast to DINAMO-S’ results presented in Figure 2, the width of the blue points (good runs) distribution is smaller, showing more robust and precise reference building. Figure 9 shows two examples of references and their uncertainties produced by DINAMO-ML and the corresponding correctly classified runs, where the left plot is a good run and the right plot depicts a bad run.
Figure 9: Top row: examples of a correctly-classified good (left, in blue) and bad (right, in red) runs by the DINAMO-ML algorithm. The corresponding references and their uncertainties are depicted by the black line and the grey shape, respectively. Bottom row represents the pull values.
For DINAMO-ML, we use the same procedure to estimate how good a model is able to capture the data variability as we used for DINAMO-S. After the z-score transformation of every histogram in the dataset and every corresponding modeled ($\hat{\bm{\mu}},\hat{\bm{\sigma}}$), we compare the resulting areas formed by standard deviations using the Jaccard distance metric. Figure 10 shows the comparison between the real data variablity and the modeled one in the transformed space and the corresponding value of the Jaccard distance. Almost perfectly overlapping shapes confirms precise model’s outputs.
Figure 10: Comparison of the references with their uncertainties produced by the DINAMO-ML algorithm and the distribution of good run histograms (where both shown in the z-score transformed space). As a measure between the two areas, the Jaccard distance is adopted.
Appendix D Models Comparison: Sensitivity and Specificity, Jaccard distance
Figure 11 compares the models in terms of sensitivity (true positive rate) and specificity (false positive rate), aggregated over 1,000 datasets. While the DINAMO-ML model outperforms the statistical-based DINAMO-S model in both metrics, the improvement is particularly evident in specificity. In other words, although both models effectively detect anomalies, DINAMO-S more frequently misclassifies good runs as bad.
Regarding the models’ ability to capture data uncertainty, Figure 12 presents the distribution of the Jaccard distance metric across 1,000 datasets, used to assess the quality of uncertainty modeling. While the median values for both models are similar (with DINAMO-ML performing slightly better), the confidence interval for the machine learning-based method is narrower, indicating more consistent uncertainty estimates.

### v2

@@ Appendix B DINAMO-S algorithm details
Algorithm 2 DINAMO-S: EWMA-based algorithm
@@ Appendix C DINAMO-ML algorithm details
This appendix presents some additional information related to the machine learning-based DINAMO model.  Figure 12 shows the logarithm of the reduced $\chi^{2}$ anomaly score for each run. Similarly to the DINAMO-S presented on Figure 4, DINAMO-ML has proven to be highly accurate in detecting bad runs as well as to quickly adapt to changing conditions.
Figure 12: Anomaly score of the DINAMO-ML algorithms as a function of the run number for a single dataset composed of 5000 runs. Red points represent the ground-truth bad runs and the blue points represent the good runs. The black dashed line is the optimized, using the historical regime, threshold to assign a class label: if a run is above it, the run is predicted to be bad and vice versa.
Algorithm 3 shows the full pseudocode of the DINAMO-ML method. It details the initialization process, the context-building procedure, the inference steps for anomaly detection, and the training mechanism that allows the model to adapt over time as new good runs are observed.
Input: $\{\bm{x}_{i},y_{i}\}_{i=1}^{N}$, buffer size $M$, batch size $K$, learning rate, etc.
Initialize:
  $\mathcal{D}\leftarrow\varnothing$
  $\rm{NNParams}\leftarrow\textsc{InitTransformerEncoder}()$
  $\rm{Optimizer}\leftarrow\textsc{AdamW}(\rm{NNParams},\rm{lr}=5\times 10^{-4},\rm{weight\_decay}=10^{-4})$
  EarlyStopPatience $\leftarrow 5$
for $i\leftarrow 1$ to NN do
  // 1. Normalize the new histogram
  Normalize $\bm{x}_{i}$ (if nonzero), get $\bm{\tilde{x}}_{i}$
  // 2. Perform inference to obtain reference prediction
  $\rm{context}\leftarrow\textsc{BuildContext}(\mathcal{D},\,i,\,M)$
  $(\hat{\mu}_{i},\hat{\sigma}_{i})\leftarrow\textsc{TransformerEncoderForward}(\rm{context},\,\rm{NNParams})$
  // 3. Compute $\chi^{2}_{\nu}$ and compare to anomaly threshold to classify
  // 4. A posteriori label is assigned (by shifter or ground truth)
  if $y^{(i)}=0$ then
   $\mathcal{D}\leftarrow\mathcal{D}\cup(\bm{\tilde{x}}_{i},i)$
   // 5. Training step
   Select the $K$ most recent entries from $\mathcal{D}$ (if fewer than $K$ exist, take all).
   $nll_{\rm{batch}}\leftarrow 0$
   Optimizer.zero_grad()
   for each $(\bm{\tilde{x}}_{k},i_{k})$ in the selected batch:
    $\rm{context}\leftarrow\textsc{BuildContext}(\mathcal{D},\,i_{k},\,M)$
    $(\hat{\bm{\mu}},\hat{\bm{\sigma}})\leftarrow\textsc{TransformerEncoderForward}(\rm{context},\,\rm{NNParams})$
@@
    $nll_{\rm{batch}}\leftarrow nll_{\rm{batch}}+nll_{\mathrm{k}}$
   $nll_{\rm{batch}}\leftarrow nll_{\rm{batch}}\;\big/\;\#(\rm{selectedbatch})$
   Backprop on $nll_{\rm{batch}}$
   CheckEarlyStopping(EarlyStopPatience)
  end if
  /* End of iteration for run $i$ */
end for
Algorithm 3 DINAMO-ML: Transformer encoder-based algorithm